### v1

# NeuCo-Bench: A Novel Benchmark Framework for Neural Embeddings in Earth Observation

Abstract
We introduce NeuCo-Bench, a novel benchmark framework for evaluating (lossy) neural compression and representation learning in the context of Earth Observation (EO). Our approach builds on fixed-size embeddings that act as compact, task-agnostic representations applicable to a broad range of downstream tasks. NeuCo-Bench comprises three core components: (i) an evaluation pipeline built around reusable embeddings, (ii) a new challenge mode with a hidden-task leaderboard designed to mitigate pretraining bias, and (iii) a scoring system that balances accuracy and stability. To support reproducibility, we release SSL4EO-S12-downstream, a curated multispectral, multitemporal EO dataset. We present initial results from a public challenge at the 2025 CVPR EARTHVISION workshop and conduct ablations with state-of-the-art foundation models. NeuCo-Bench provides a first step towards community-driven, standardized evaluation of neural embeddings for EO and beyond.
Cartoon summary of the NeuCo-Bench framework. Multi-temporal, multi-modal, and multi-channel inputs $x$ are compressed into fixed-size embeddings $z=E(x)$ by a user-defined encoder $E$. The embeddings are then linearly probed on a diverse set of regression and classification downstream tasks to assess the general-purpose quality of $z$.
1 Introduction
The rapid growth of visual data, from online media to scientific observation, has made efficient compression a central challenge for storage, transmission, and large-scale analysis (Pouyanfar et al., 2018; Wang et al., 2018a; Gomes et al., 2025). Traditional codecs such as JPEG2000 (Skodras et al., 2001) and more recent learned autoencoders (Ballé et al., 2016) are optimized for pixel-level distortion, largely reflecting human visual perception. However, many machine learning pipelines care less about perceptual fidelity and more about semantic fidelity, retaining the information needed to solve downstream tasks (Huang and Wu, 2024). This gap is particularly critical in domains like Earth Observation (EO), where petabyte-scale datasets of multi-modal satellite imagery must support diverse analytical tasks ranging from environmental monitoring to disaster response (Guo et al., 2017). EO data are characterized by substantial redundancy and noise across multiple spectral bands and temporal sequences, amplifying the need for compression strategies that efficiently capture underlying, task-relevant information (Gomes et al., 2025). This gives rise to the question: How much task‐relevant information can be squeezed into compact data representations?
Recent work has shown that compressed latent representations can preserve rich semantic content, enabling pipelines to operate directly on features without reconstructing the input image (Torfason et al., 2018; Singh et al., 2020). Self-supervised foundation models (FMs) further demonstrate that embeddings can transfer across tasks with minimal fine-tuning. Yet, their dimensionality often rivals or exceeds the size of the original data, reintroducing storage and bandwidth bottlenecks (Gomes and Brunschwiler, 2024; Lu et al., 2024). Despite these advances, there is currently no standardized framework evaluating how effectively compressed representations retain semantic content across multiple downstream tasks. Existing evaluations remain fragmented, often restricted to pixel fidelity, single-task utility, or unconstrained high-dimensional embeddings, making it challenging to compare approaches on a common basis.
To address this, we introduce NeuCo-Bench, a model-agnostic benchmark for assessing the semantic quality of embeddings in EO. Our frameworks is designed to (1) evaluate compressed embeddings under strict size constraints, (2) probe semantic retention using linear models across diverse downstream tasks, (3) support multi-modal and multi-temporal data typical of data-intensive EO settings, and (4) foster community contributions, including new datasets and compressors—towards establishing open, task-centric compression standards. Our key contributions are:
  • Section˜3Benchmarking Framework: We develop a standardized framework for evaluating compressed embeddings via downstream tasks, aligning with task-centric machine‐to‐machine workflows.
  • Section˜4Benchmark Tasks: We curate and release a suite of novel EO downstream tasks, spanning cloud analysis, agricultural monitoring, forest quantification, urban heat islands identification, and land cover analysis.
  • Section˜5Benchmark Evaluation: We validate the utility of NeuCo-Bench through the 2025 CVPR EARTHVISION data challenge, introducing a novel hidden-task evaluation scheme. We further test embedding quality under diverse compression strategies, including pre-trained neural compressors and FMs.
2 Related Work
Below, we review research fields relevant to contextualize NeuCo-Bench:
Classical rate-distortion compression. Image and video codecs such as JPEG, JPEG2000, H.264/HEVC (Wallace, 1991; Skodras et al., 2001; Sullivan et al., 2012; Richardson, 2010) exploit handcrafted transforms (Goyal, 2001; Bracewell, 1986; Daubechies, 1992) and entropy coding to reduce statistical redundancy. Their performance is evaluated through the rate–distortion (RD) trade-off between compressed bit rate and reconstruction fidelity (e.g., MSE, PSNR).
Neural image compression. Learned autoencoders replace handcrafted transforms with analysis and synthesis networks jointly optimized for rate and distortion. Differentiable entropy models enable superior RD performance compared to JPEG2000 (Ballé et al., 2016; Theis et al., 2022), with subsequent extensions using hyperpriors (Ballé et al., 2018; Minnen et al., 2018), autoregressive models (Minnen and Singh, 2020), and transformers (Qian et al., 2022). With automated vision pipelines, the concept of compression for machines shifts focus from human‐perceptual fidelity to task‐driven utility. End-to-end approaches jointly optimize compressors with task networks (Chamain et al., 2020, 2021; Le et al., 2021; Codevilla et al., 2021; Wang et al., 2021, 2023a; Fischer et al., 2025). Other methods enforce invariance to task-relevant augmentations through self-supervised objectives (Dubois et al., 2022), or bypass reconstruction by training tasks directly on compressed latents (Torfason et al., 2018; Duan et al., 2023; Singh et al., 2020).
Compression in EO. EO imagery presents unique compression challenges, with multi-spectral bands, temporal sequences, and petabyte-scale archives (Guo et al., 2017; Wilkinson et al., 2024). Traditional pipelines often rely on codecs like JPEG2000 (Yeh et al., 2005). Recent neural approaches extend rate-distortion autoencoders to EO imagery, achieving significant rate-distortion improvements on multispectral data (Alves de Oliveira et al., 2021; Kong et al., 2021; Cao et al., 2022), while temporal compression remains underexplored (Du et al., 2024; Wang et al., 2018b). For a comprehensive review, see (Gomes et al., 2025). Importantly, most works evaluate RD, not task relevance.
EO Foundation Models. Self-supervised learning has enabled large-scale vision foundation models (FMs) pretrained on vast, unlabeled satellite datasets using masked reconstruction, contrastive, or predictive tasks (Wang et al., 2022a; Sun et al., 2022; Wang et al., 2022b; Mai et al., 2022; Wang et al., 2023b; Hong et al., 2023; Jakubik et al., 2023; Liu et al., 2024). These FMs produce versatile high-dimensional embeddings for EO downstream applications, such as flood segmentation, land-use mapping, and environmental monitoring. More recently, multimodal EO foundation models have begun fusing data modalities, such as SAR and optical imagery, to capture diverse geophysical characteristics and improve application performance (Li et al., 2022; Fuller et al., 2023; Xiong et al., 2024; Jakubik et al., 2025; Brown et al., 2025). However, with the exception of (Brown et al., 2025), the resulting latent representations often rival or exceed the original data size, creating data transfer and data processing bottlenecks. Gomes and Brunschwiler (2024) addresses these challenges by integrating neural compression into FM bottlenecks. On image level, Rolf et al. (2021) utilizes fixed, random convolutional kernels to engineer features as basis for linear regression to predict user labels.
EO benchmarks. Current EO domain benchmarks, such as GEO-Bench (Lacoste et al., 2023) and PANGAEA (Marsocci et al., 2025), evaluate FMs by fine-tuning backbones or training complex decoders on intermediate features. These approaches typically require model access and significant computational resources, with limited consideration given to factors such as embedding size and workflow efficiency. In contrast, NeuCo-Bench evaluates fixed-size embeddings through task-agnostic linear probing without any need for access to model backbones. In fact, our approach treats the encoder as a black box that converts any input to a given number of features. NeuCo-Bench provides a lightweight, size-aware evaluation protocol for efficient local testing, and it is structured as a flexible, extendable framework, designed to accommodate future downstream tasks and evaluation methods. Moreover, NeuCo-Bench can be deployed as a novel challenge format that simulates real-world scenarios by requiring participants to submit compressed EO embeddings without prior knowledge of the specific downstream tasks. This setup reflects the demand for broadly generalizable embeddings. As demonstrated in the 2025 CVPR EARTHVISION data challenge, NeuCo-Bench integrates with established platforms such as EvalAI (Yadav et al., ), and is designed to support future competitions on new, unseen tasks.
3 Benchmarking Framework
Figure 1: Workflow diagram: The user of our benchmark compresses a set of downstream data into fixed-length embeddings (size $N=1024$ here). Our benchmark loads the embeddings for each of the $T$ downstream tasks, performs $K$ evaluations, randomly sampling a training and test split from the task data each evaluation, and scores the result compared to all previous experiments as stored in the Scoring Database.
At the heart of the NeuCo-Bench framework resides (i) an embedding evaluation workflow and (ii) a ranking method to fairly compare performance across multiple tasks of varying difficulty.
Evaluation workflow. Figure˜1 visualizes the pipeline for an Experiment $p$ compressing the samples indexed by $i$ of Benchmark data to create a set of fixed-size embeddings (Embedding set) through an Encoder $E^{(p)}$: These embeddings $X^{(p)}$ are provided to NeuCo-Bench, which performs the evaluation given corresponding Task labels (aka downstream tasks $t=1\dots T$) to return the Benchmark results through a Leaderboard. For each Experiment, the NeuCo-Bench framework performs an Embedding Evaluation given Multimodal, multitemporal datacubes across a set of downstream tasks undisclosed to the developers of a given Compression Model comprising an Experiment $p$. Correspondingly, our framework aggregates scores $s_{t,k}$ per training and test split $k$ to gather statistics for the quality score $Q_{t}$ per downstream task $t$. Consequently, a Scoring algorithm applies a task difficulty-dependent ranking scheme. A standalone Python implementation is available online [^1] https://github.com/embed2scale/NeuCo-Bench. Section˜B.2 provides additional details.
Evaluating embeddings. For NeuCo-Bench, each input sample must be represented as a fixed-size embedding to compress an input data cube, e.g., our EO downstream tasks as detailed in Section˜4, or future extensions, cf. Section˜6. Currently, image-level linear regression and binary (softmax) classification are supported. NeuCo-Bench enforces embeddings of fixed, but configurable, size but otherwise does not constrain how embeddings are generated. Following Fig.˜1, NeuCo-Bench evaluates the compressed embeddings $X^{(p)}$ of an experiment $p$ as follows: For each task $t=1\ldots T$, $K$ linear classifiers (with $N$ tunable parameters $a_{1\dots N}$ plus bias term $a_{0}$) are trained to fit the downstream task labels $y_{\text{ID}}$. Each $k=1\dots K$ denotes a separate, randomly generated split of the downstream task $t$ into a training and testing set. For each tuple $(t,k)$ NeuCo-Bench computes an accuracy measure $s_{t,k}$, utilizing $R^{2}$ (R-squared) for regression tasks and the $F_{1}$ score for (binary) classification. From the set $\left\{s_{t,k}\right\}_{k=1\dots K}$, NeuCo-Bench derives a signal-to-noise-like quality score $Q_{t}^{(p)}$ as the mean performance on task $t$ sensitive to the variability in performance of experiment $p$:
$Q_{t}^{(p)}=100\epsilon\frac{\left\langle s_{t,k}\right\rangle_{k}}{\text{std}_{k}\left(s_{t,k}\right)+\epsilon}.$ (1)
Here, $\left\langle\cdot\right\rangle_{k}$ denotes averaging and $\text{std}_{k}\left(\cdot\right)$ the standard deviation as calculated over the $K$ splits. The parameter $\epsilon>0$ acts as a regulator avoiding high variability in $Q_{t}^{(p)}$ for small $\text{std}_{k}\left(s_{t,k}\right)$. The quality score Eq.˜1 varies in $[0,100]$ for both, classification and regression. Thus, $Q_{t}^{(p)}$ allows for an interpretation of mean accuracy in percent. Compared to using the mean $R^{2}$ over the $K$ splits, $Q_{t}^{(p)}$ penalizes methods with larger variance in the $R^{2}$. Further details on the quality score is provided in Section˜A.1 of the suplpementary material. We note: experiments that perform worse than simply predicting the mean of labels $y_{\text{ID}}$ for regression tasks result in negative $s_{t,k}$ degrading the mean performance. In fact, compression models with negative $Q_{t}^{(p)}$ should be flagged unreliable—they seriously underperform.
Task difficulty-dependent ranking. A novel scoring method is introduced by NeuCo-Bench. It is designed to compare the overall performance of multiple participants over multiple tasks. Based on a rank-then-aggregate approach (Wiesenfarth et al., 2021), NeuCo-Bench dynamically weights the performance across tasks depending on their relative difficulty: Each experiment initially receives a rank $R_{t}^{(p)}$ per task, with the best rank given to the experiment with highest $Q_{t}^{(p)}$. To break ties, all tied experiments are given the lower (better) rank. An experiment’s final rank is calculated from the weighted mean rank across all tasks:
$s^{(p)}=\sum_{t=1}^{T}w_{t}R_{t}^{(p)}\quad\text{with}\quad w_{t}=\text{std}_{p}\left(Q_{t}^{(p)}\right)\Big/\sum_{t=1}^{T}\text{std}_{p}\left(Q_{t}^{(p)}\right)$ (2)
where the tasks are weighted by the standard deviation of the $Q_{t}^{(p)}$ of all experiments on the task. The weighting scales the importance of the tasks such that (a) tasks where all participants perform similarly receive low importance, and (b) tasks where the participants differentiate between each other are weighted highly. NeuCo-Bench also provides the mean $Q$ value $\langle Q_{t}^{(p)}\rangle_{t}$ as an experiment-specific measure of performance. For scenarios with few experiments where the interpretation of a ranking is limited in terms of task difficulty, $\langle Q_{t}^{(p)}\rangle_{t}$ serves as an alternative metric to compare (individual) experiments. Section˜A.1 provides additional analysis of the ranking scheme.
4 Benchmark Tasks
NeuCo-Bench provides a set of pre-processed, heterogeneous downstream tasks designed for continuous extension in the future. The initial release provides regression labels that are easily turned into binary classification tasks through a threshold.
Table 1: Summary of spatial coverage for current set of data cubes and associated number of downstream tasks.
Dataset
Spatial
Coverage
Temporal
Coverage
Years
of Labels
# Samples
# Tasks
[…]
Landcover Europe 2018 2018 4691 2
Biomass Global 2019 2019 2415 2
Clouds Global 2018 - 2020 2018 - 2020 1140 1
Heatisland Northern Hemisphere 2022 2021 - 2024 1659 2
Figure 2: Spatial distribution of the downstream tasks.
The Earth observation satellite data of the downstream tasks summarize in Table˜1 follow the SSL4EO-S12 data cube structure (Blumenstiel et al., 2025; Wang et al., 2023b). We utilize 13 channels of Sentinel-2 Level-1C Top-of-the-Atmosphere (S2L1C) and 12 channels of multi-spectral Sentinel-2 Level-2A surface reflectance (S2L2A). On top we spatially align 2 channels of radar Sentinel-1 (S1) GRD product polarizations (VV and VH). For a given geolocation we retrieve four timestamps, one per season: winter (Dec–Feb), spring (Mar–May), summer (Jun–Aug), and fall (Sep–Nov). Figure˜1 depicts these four seasonal data cubes, with each containing 27 bands. Google Earth Engine (Gorelick et al., 2017; GEE, ) (GEE) was utilized to download all relevant satellite data. All labels, except for the Clouds use case, have been retrieved from GEE, too. Following (Blumenstiel et al., 2025), the processed data was stored as cloud-ready ZIP-Store of the Zarr file format.
The downstream tasks contain between 1100 and 4691 samples (locations/labels), which are distributed globally (Fig.˜2). The associated satellite data cubes are pre-processed and filtered to ensure UTM-projected patches with a size of 264 × 264 pixels without spatial overlap.
The crops task covers cropland in the US Corn Belt and is provided by the US Department of Agriculture (USDA) (Boryan et al., 2011). Soybean and corn were selected as primary focus classes for this downstream task, with the fraction of corn and soybean within each patch serving as the label. The Crop Data Layer is published annually with a spatial resolution of 30 meters. The labels are available as post-processed data222e.g., the label year 2023 corresponds to crops cultivated during the 2022 growing season.
The landcover tasks leverages aggregated land use data from the European Environment Agency (EEA), which includes various land cover classes such as forests, urban areas, water areas, and agricultural land (European Environment Agency (2018), EEA). These labels represent the dominant land cover within each patch at a spatial resolution of 100 meters within Europe. Based on this data, two downstream tasks are provided for forests and agricultural land 2018 within the challenge.
The biomass tasks uses above-ground biomass estimates derived from LIDAR measurements from the Global Ecosystem Dynamics Investigation (GEDI) instrument. GEDI provides structural information on vegetation height and density, allowing robust models to estimate above-ground biomass in megagrams per hectare (Mg/ha) (Dubayah et al., 2022). Within NeuCo-Bench, the GEDI Level 4A biomass estimates were spatially aggregated to the satellite patches with 264x264 pixels, providing a mean biomass value and its standard deviation as regression targets.
The clouds provides cloud cover fractions based on CloudSen12+ (Aybar et al., 2024) as labels and pre-processed Sentinel-1 and Sentinel-2 data cubes as corresponding observations. Although the SAR data is not affected by clouds, Sentinel-1 is included alongside Sentinel-2 to ensure a consistent data structure for all downstream tasks.
For the heatisland use case, Landsat-8 Land Surface Temperature (LST) provides surface temperature data that are used as labels for urban areas (Observation and Center, 2020). This is particularly relevant in the context of heat events and future urban planning, and contains 1659 samples. The corresponding tasks address the mean surface temperature and its standard deviation per data cube stack. Further information on the data and their downstream tasks during the competition can be found in the Section˜B.1.
5 Benchmark Evaluation
We describe how we tested NeuCo-Bench in a real‐world setting via CVPR EARTHVISION 2025, by detailing the experimental setup in Section 5.1, as well as discussion of outcomes and learnings in Section 5.2. We further present baseline evaluations exploring our set of downstream tasks in Section 5.3.
5.1 Data Challenge Validation
To validate NeuCo-Bench under realistic conditions, we utilized it in the 2025 CVPR EARTHVISION data challenge. Participants were tasked with compressing multi‐modal, multi‐temporal EO imagery, cf. Section˜4, into 1,024-dimensional embeddings. Given the SSL4EO-S12 input data cubes, this amounts for a compression ratio of approx. 7,000. Crucially, participants did not know which or the number of downstream tasks their embeddings would be evaluated on; this hidden‐task design discourages overfitting and encourages the development of general‐purpose EO representations. Participants were ranked, according to Section˜3, across two sets of downstream tasks. One modification was made to the dataset compared to the dataset described in Section˜4; the clouds task targets were mainly replaced by zeros, causing heavy skewness in the labels and basically random connection between the imagery and labels.
Phases. In a three‐week development phase, teams developed embedding methods using the publicly available SSL4EO v1.1 dataset. A partial release of 5 tasks, each with a subset of its samples but no information of the task type, allowed participants to receive initial feedback for development. Submissions returned only the mean $Q$ value to prevent leakage of task‐specific performance and information.
In the subsequent three‐day evaluation phase, an extended set of 9 tasks, and new data on the tasks also used in the development phase, was released. The teams had three days and up to three submissions to encode and submit embeddings; these runs defined the final leaderboard standings. By the end, two winning teams were chosen; The first based on the dynamic ranking scheme, and the other as the team with highest mean $Q$ score.
Platform and infrastructure. The NeuCo-Bench framework was modified to utilize Eval.AI (Yadav et al., ) to collect submissions. NeuCo-Bench ran on a separate 8-vCPU server, retrieved new submissions via API, executed the evaluation, and pushed results to a custom leaderboard hosted on GitHub, displaying the dynamic ranking described in Section˜3, as well as back to Eval.AI. Additional details are available in Section˜A.2.
5.2 Data Challenge Results
Participation and ranking. Twenty‐three teams submitted to the development phase; sixteen went on to the final evaluation, nine of which shared their submissions publicly. The quality scores $Q_{t}^{(p)}$, shown in Fig.˜3 for the evaluation phase, varied widely from 0–5 on some tasks to 5–40 on others. The evaluation method efficiently scaled task importance ensuring that the tasks impacted the leaderboard relative to their differentiating effect across the teams. Notably, the weighting reduced the impact of the tasks with zero/random labels. Further, the dynamic ranking caused a swap between the original first and second place methods due to a third team, highlighting the impact of adaptive weighting.
Figure 3: Quality score $Q_{t}^{(p)}$ of the participants of the data challenge evaluation phase with corresponding task weight used for ranking. Teams are ordered by their final leaderboard rank, with the winner as Team 1. Team 10 is a simple averaging baseline described in Section˜5.3 and 11 denotes a baseline of normally distributed random embeddings.
Top methods. The team that achieved the best overall rank and the team with the highest average $Q$-Score both built their embeddings by ensembling multiple FM representations, although with different approaches; One pre-training backbones, and the other training a bottleneck based on frozen FM backbones. The fourth-place team took a different path, forgoing any pre-training and instead generating embeddings using the MOSAIKS method (Rolf et al., 2021).
Key takeaways. The Embed2Scale data challenge at the CVPR EARTHVISION 2025 demonstrated that NeuCo-Bench efficiently evaluates and ranks the performance of compact embeddings over multiple downstream tasks. The scoring method produced a more balanced and discriminative ranking compared to uniformly weighting the tasks, particularly noticeable in the down-weighting of the two random tasks. Hiding the tasks efficiently prevented overfitting, ensuring fairness between the participants. Both winning solutions were based on FMs, indicating these can indeed provide semantically rich, general embeddings. However, also non-FM based solutions scored high.
5.3 General Evaluations
We assess the NeuCo-Bench framework through a series of experiments spanning three perspectives. First, we follow the challenge setup by constraining all methods to produce 1,024-dimensional embeddings from the input EO data cubes (see Section˜4). We introduce and test approaches including (i) learned neural compressors optimized for rate–distortion trade‐offs, and (ii) self‐supervised FMs, spanning both uni- and multimodal architectures. Second, we relax this requirement to analyze the impact of embedding dimensionality on downstream performance. By varying the size of CNN- and ViT-based embeddings, we investigate whether larger or smaller representations alter relative performance, motivating the 1,024-dimension default and demonstrating the flexibility of the benchmark for exploring embedding size as an additional axis of evaluation. Lastly, we revisit and explore the assumption of linear probes as decoder models.
Data and assumptions. Each input consists of four seasonal snapshots from Sentinel-1 (radar) and Sentinel-2 (optical L1C and L2A). To meet the embedding‐size constraint, we employ frozen, pretrained backbones followed by downsampling, and adopt simplifying assumptions: (1) although S1 and S2 capture complementary signals, substantial correlation justifies unimodal baselines, reflecting common practice in neural compression and FM development on EO data; (2) We fuse the four seasonal views into a single tensor, either before encoding (pre‐encoding aggregation) or after encoding (post‐encoding aggregation), to produce one 1,024‐dimensional vector per sample.
Method Pretraining Loss Backbone Input Downsampling Original latent dim.
TerraMind (FM) TerraMesh Cross-Entropy Multi-modal ViT-B/16 All (S1,S2) Patch averaging $196\times 768$
MAE ViT (FM) SSL4EO MAE ViT-B/16 S2L1C Patch averaging $196\times 768$
DINO ViT (FM) SSL4EO DINO ViT-B/16 S2L1C Patch averaging $196\times 768$
DINO ResNet (FM) SSL4EO DINO ResNet-50 S2L1C Global avg., pairwise channel mean $2048\times 7\times 7$
Factorized Prior SSL4EO Rate-Distortion CNN S2L1C Adaptive pool ($4\times 4$), pairwise channel mean $128\times 14\times 14$
Averaging Baseline All (S1,S2) bilinear avg., flatten $4\times 4\times 8\times 8$
Figure 4: (Left) Radar plot of per‐task $R^{2}$ performance for a rep. subset of embedding–compression methods. Each axis corresponds to one downstream prediction task; the center denotes $R^{2}=0$, negative values are clipped to zero for clarity. (Right) Summary of method configurations.
Methods and baselines. We include a simple averaging baseline, which applies spatial downsampling, channel‐wise averaging, and flattening to yield a 1,024‐dimensional vector; this serves as a minimal reference point. Next, our neural rate–distortion compressors are implemented via Factorized Prior autoencoders (Ballé et al., 2016) pretrained on SSL4EO S2L1C data. We extract latent bottleneck features before entropy coding (cf. (Torfason et al., 2018)) and subsequently downsample them to conform to the 1,024‐dimensional limit. Finally, we evaluate several FMs based on ResNet (He et al., 2015) and ViT (Dosovitskiy et al., 2020) backbones pretrained on SSL4EO data cubes with masked‐autoencoding (MAE (He et al., 2021)) or contrastive (DINO (Caron et al., 2021)) objectives, as well as multimodal architectures like TerraMind (Jakubik et al., 2025), which are pretrained jointly on radar and optical inputs.
Results and discussion. Figure˜4 presents the linear‐probing performance in terms of $R^{2}$ for all downstream tasks assembled and a representative subset of compression methods. Downsampled embeddings from our neural rate–distortion compressors outperform the simple averaging baseline but remain below $R^{2}=0.5$. This observation highlights the characteristics of our setup with high compression rate $\sim$7,000 while testing against linear‐probe. FM embeddings reveal a task‐dependent trend: FMs in general, with contrastive (DINO) and multimodal models (TerraMind) in particular, achieve high $R^{2}$ on semantic tasks where multi-pixel context is relevant (e.g., land‐cover proportion). However, certain FMs struggle on geophysical predictions of quantities resolved at the sub-pixel level (e.g., biomass estimation), whereas multimodal and MAE FMs strike a better balance across tasks.
Temporal aggregation. Across all methods, post‐encoding aggregation consistently outperforms pre‐encoding aggregation. Accordingly, the results in Figure˜4 employ post‐encoding aggregation exclusively, despite incurring a $\times 4$ increase in embedding‐generation runtime. The performance gains are modest for static‐feature tasks (e.g., land‐cover), yet substantial for temporally sensitive tasks (e.g., cloud‐fraction estimation), underscoring the importance of preserving per‐snapshot details.
Embedding size. In Fig.˜5(a) we numerically study how the embedding size impacts performance testing the models introduced above.
  • CNN backbones. Performance peaks for embedding sizes $128\lesssim N\lesssim 1024$, with accuracy dropping outside. Larger embeddings add computational cost without gain in performance.
  • ViT backbones. We find the best performance at $N=1024$, the upper limit allowed by the embedding dimension. A lower $N$ consistently reduce accuracy—except for certain regression tasks such as for Biomass.
  • Trade-offs. While larger embeddings increase the number $N=|A|$ of (linear) probe parameters $A$ (cf. Fig.˜1), smaller $N$ often fail to retain task-relevant semantics.
Our experiments support an embedding size of $N=1024$ as a balanced default across downstream tasks, cf. Fig.˜2. The NeuCo-Bench framework is flexible to explore size–utility trade-offs, analogous to rate–distortion analysis in neural compression. Section˜B.3.3 provides additional and detailed per-task performance plots.
((a))
((b))
Figure 5: (a) Average downstream $R^{2}$ as a function of embedding size $\log N$. Largest size is the full channel count (CNN) or native patch-token dimension (ViT). (b) Average downstream $R^{2}$ comparing linear probing against one-layer and two-layer MLP probes on 1024-dimensional embeddings.
Linear probing assumption. We evaluate embedding quality using linear probing, a widely adopted practice in representation learning (Xu and Tewari, 2021) to focus on embeddings without fine-tuning encoder backbones. While non-linear probing (e.g., small MLP heads) can in principle capture richer structures, it risks compensating for poor embedding quality Plachouras et al. (2025). Our experiments in Fig.˜5(b) and Section˜B.3.3 demonstrate: Replacing linear probing by small, non-linear decoders yields only marginal gains for top-performing embeddings, while providing larger improvements for weaker ones. Further, non-linear probing substantially increases computational cost. Thus, linear probing remains an (energy-)efficient and reliable measure of how much semantically relevant information is directly accessible from an embedding space. Linear probing efficiency enabled NeuCo-Bench at the 2025 CVPR EARTHVISION data challenge to run over 400 submissions within minutes on commodity hardware.
6 Future Work
Reproducibility. Data contributions to NeuCo-Bench demands a permissive CC-BY 4.0 license. Our 2025 CVPR EARTHVISION data challenge required the winners to release their solution under Apache 2.0 license [^3] https://creativecommons.org/licenses/by/4.0 and https://www.apache.org/licenses/LICENSE-2.0.html. The same holds for any future extension of NeuCo-Bench where Section˜B.5 provides further details. While Section˜A.1 has background on the theoretical basis of our evaluation metric with references to code, Sections˜A.2 and B.4 share specifics on the hardware, the software environment, and hyperparameter settings NeuCo-Bench runs on. Section˜A.3 provides some additional elements of running the 2025 CVPR EARTHVISION data challenge. Sections˜B.1 and B.2 gathers general facts on the benchmark data and code framework, respectively. Section˜B.3.2 is dedicated to additional model performance metrics for downstream tasks provided by NeuCo-Bench.
Fixed-size compression. Our current evaluation emphasizes fixed-size embeddings, as fixed-size vectors enable fast retrieval, comparison, and inference, critical for machine-oriented downstream tasks. Nevertheless, the framework can be naturally extended to incorporate entropy coding, where embeddings are further losslessly compressed for transmission before decoding and use. In this setting, the proposed performance scores directly evaluate utility as a function of the final entropy-coded bitrate, thereby bridging task performance with classical rate–distortion analysis.
Choice of tasks. The discriminative power of results depends on downstream tasks. We curated diverse, image-level tasks focused on global semantic content, together with a dynamic ranking scheme. Future extensions will include spatially structured tasks such as pixel-level segmentation or time-sensitive predictions, which may require less aggressive compression ratios than the current value of $\sim$7,000. Probing strategies may also evolve as tasks grow in complexity.
Downstream data. Although the current benchmark is rooted in EO, its design is domain-agnostic. Extensions can cover multi-modal, spatio-temporal data across domains such as weather forecasting, medical imaging, or autonomous driving. Our experiments leveraged SSL4EO-S12 as an initial sweet spot (multi-modal, multi-temporal, multi-spectral). The same concept readily transfers along the axes of domain, modality, time, and channels.
Building a NeuCo-Bench community. With NeuCo-Bench, we provide a seed to grow an eco-system centered around benchmarking highly-compressed embeddings on a set of standardized, community-contributed downstream tasks. To avoid contributing to an ever-growing number of benchmark datasets, we intend to harmonize with existing ones, such as GEO-Bench and PANGAEA. NeuCo-Bench’s mission is to provide a standardized framework of benchmarking embeddings. We warmly welcome contributions from all research areas involved in neural compression. Future extensions of the benchmark framework include pixel-wise and temporal downstream tasks.
Societal impact & information privacy. NeuCo-Bench fosters open science for neural compression. SSL4EO-S12 derives from the Copernicus Programme under a free and open data policy. For label sources, please check details in Section˜B.5. The provided dataset is void of detailed time and geolocation information, sparsely spread in time with a single timestamp per quarter, and limited to 10m pixel resolution, limiting any surveillance applications. For benchmarks with strong compression ratio, NeuCo-Bench may act as a privacy conserving approach where pixel-level reconstruction becomes impossible. NeuCo-Bench does not touch on any concerns related to human rights. We include the Linux Foundations Code of Conduct for future contributions to the code released with this paper.
7 Conclusion
We presented NeuCo-Bench, the first of its kind, task-driven benchmarking framework for compression that evaluates neural embeddings by downstream task performance, rather than pixel fidelity. The framework introduces a novel rank-then-aggregate scoring method which dynamically determines the task complexity based on score statistics. We demonstrated the NeuCo-Bench framework by setting up the CVPR EARTHVISION 2025 data challenge on multi-modal, multi-temporal, and open-source EO data. We introduced a novel set of real-world downstream tasks which remained undisclosed at the time of the data challenge, and have been publicly released after the conclusion of the competition. NeuCo-Bench encourages the development of methods that generate semantically-rich, general-purpose embeddings.
For our setup, experiments demonstrated that multi-modal foundation models yields strong overall performance—particularly on semantic land-cover tasks. Post-encoding fusion of seasonal views resulted in notable gains for temporally sensitive tasks such as cloud cover prediction. We also observed that smaller and, in some cases, larger embedding sizes degrade performance. This observation highlights compact embeddings as a practical choice for image-level tasks when high-quality annotations and compute resource become scarce.
NeuCo-Bench is open source and ready for extension—either by novel evaluation methods or additional downstream tasks without any conceptual restriction to Earth observation. Currently, the NeuCo-Bench framework is limited to image-level tasks, but future work aims to extend NeuCo-Bench’s functionalities to include pixel-wise outputs, options beyond linear probing, and an assessments of bit-rate efficiency.
Acknowledgment
This work was funded by Embed2Scale which is co-funded by the EU Horizon Europe program under Grant Agreement No. 101131841. Additional funding for this project has been provided by the Swiss State Secretariat for Education, Research and Innovation (SERI) and UK Research and Innovation (UKRI).
We thank Tim Reichelt (Oxford U) and Damien Robert (U Zurich) for support in creating labels for the downstream tasks. Johannes Jakubik and Benedikt Blumenstiel (IBM Research) provided valuable discussions when writing this article. CMA is grateful for Columbia University’s academic hospitality in 2025.
Moreover, we are grateful for the active participation in our CVPR EARTHVISION 2025 workshop data challenge with valuable Q&A in https://github.com/DLR-MF-DAS/embed2scale-challenge-supplement/issues?q=is%3Aissue. In particular, conversations with Burak Ekim (UniBW) and Isaac Corley (Wherobots) sparked lively discussions, and our meetings with the winning teams KTH and Friends (KTH Royal Institute of Technology, Sweden) and 404 Embedding Not Found (Microsoft Research, USA).
Appendix A Technical Details of the Data Challenge
Before we delve into generic considerations regarding NeuCo-Bench in Appendix˜B, we introduce its origin spared by the innocent question:
If Geospatial Foundation Models claim to generate informative, generic feature vectors for a broad range of use cases, why can’t we put that claim to the test in a data challenge at this year’s CVPR? Catch: We will not disclose the downstream tasks, but simply ask to embed /compress Earth observation data.
A.1 Challenge Evaluation Method and Configuration
An individual, local score of embedding quality. As detailed in the main article, the central evaluation metric serving as quality score to answer the question above reads like
$Q_{t}^{(p)}=100\epsilon\frac{\left\langle s_{t,k}\right\rangle_{k}}{\text{std}_{k}\left(s_{t,k}\right)+\epsilon}\equiv Q=100\epsilon\frac{\bar{s}}{\Delta s+\epsilon}\quad\text{with}\quad\epsilon=0.02\quad.$ (3)
For fixed $\epsilon>0$, the maximum value of $Q$ reaches $100\epsilon\epsilon^{-1}\bar{s}=100\bar{s}$ when the statistical fluctuations vanish, $\Delta s\to 0$. Given $\Delta s\geq 0$ and $\bar{s}\in[0,1]$ [^4] For an $R$-squared score (regression task), $s<0$ penalizing good, positive values $s\in[0,1]$. In fact, negative ss–values indicate that the downstream task prediction is worse than a model simply predicting the value of the mean label. derived from a measure such as the F1-score or R-squared, the range of the quality score can be interpreted as a percentage of quality.
The numerical value of the regulator $\epsilon$ determines the scale at which $Q$ becomes insensitive to statistical fluctuations $\Delta s$: As long as $\Delta s\gg\epsilon$, in zero-order approximation, we have $q=Q/100\approx\bar{s}/\Delta s$ a measure of signal-over-noise for the quantity $s$. At the other end of the spectrum when $\epsilon\gg\Delta s$ dominates the noise, we conclude $q\approx(1-\Delta s/\epsilon)\bar{s}<\bar{s}$ in first order of $\Delta s/\epsilon$. However, when $\Delta s\approx\epsilon$, $q\approx\bar{s}/2$ is relatively insensitive to the noise $\Delta s$. In particular, when the score $s$ varies about $\Delta s\approx 0.02=2$$\approx\epsilon$ across the set of validations indexed by $k$, then $Q\approx 50\bar{s}$, i.e. for almost perfect $s\approx 1$ values across the board, we obtain a $Q$ close to 5050%. Only, when $\Delta s$ significantly drops below the fixed $\epsilon=2$%, $Q$-scores close to 100% are possible (given close to perfect $s$-scores).
In order to gather sufficient statistics to fairly compare the challenge participants, the number of linear classifiers trained on separately-sampled training and test sets was varied from $k=1,2,\dots,40$ during the development phase and $k=1,2,\dots,200$ during the final evaluation phase. While the seed for the random number generator used for the training and test set splits is kept constant for NeuCo-Bench in Appendix˜B, for the CVPR EARTHVISION data challenge it was initialized at random. Our choice was motivated by the effort to minimize information leakage about the hidden downstream tasks to the data challenge participants. During the development phase, submissions could test the constant set of predefined downstream tasks over a three-week period and submit 10 times a day.
Global ranking relative to other challenge participants. On top of a single participants $p$’s (local) performance score $Q_{t}^{(p)}$, we added a global ranking scheme as follows: Both local and global rankings assign rank $R_{t}^{(p)}$=1 to the highest performing participant and ascending rank $R_{t}^{(p)}$-values for decreasing performance. Ties are broken such that all tied participants get the lower (best) rank. The algorithmic design of our approach is best illustrated in a Python code implementation like:
1q = {
2 ’team1’: 13.223,
3 ...,
4 ’teamP’: -3.55677
5}
6
7def rank(q:dict, descending:bool = True) -> dict:
8 sign = 1
9 if descending:
10 sign = -1
11 return {
12 p: 1 + len(
13 [ s_sub for s_sub in q.values() if sign*s_sub < sign*s ]
14 )
15 for p, s in q.items()
16 }
17
18ranked_q = rank(q)
where the Python dictionary q serves as input to rank() to generate $R_{t}^{(p)}$=ranked_q and the boolean parameter descending triggers whether the highest or lowest value is deemed best. Utilizing rank(), the local ranking $R_{t}^{(p)}$ orders participant $p$ on task $t$ with the highest (best) score $Q_{t}^{(p)}$, descending=True. The second, global ranking across tasks assigns rank $R^{(p)}=$1 to the participant with the lowest (best) weighted average rank score
$s^{(p)}=\sum_{t=1}^{T}w_{t}R_{t}^{(p)}\quad\text{where}\quad w_{t}=\text{std}_{p}Q_{t}^{(p)}\Big/\sum_{t=1}^{T}\text{std}_{p}Q_{t}^{(p)}$ (4)
by setting descending=False. In contrast to $\text{std}_{k}$ over cross-validation folds $k$ in Eq.˜3, here, $\text{std}_{p}$ runs over the number of participants $p$ of a fixed task $t$, i.e., the weight $w_{t}$ computes the variation of our evaluation metric $Q^{(p)}_{t}$ for a given task $t$ across all data challenge participants $p$. Thus, $w_{t}$ serves as a measure of task competitiveness to characterize and automatically distinguish tasks $t$.
Our design rationale of the weighted_score for the 2025 CVPR EARTHVISION data challenge was as follows:
  • reward participants scoring well for a given downstream task
  • discount the quality score $Q$ depending on the task competitiveness of a downstream task, i.e., measure relative performance among challenge participants for a given downstream task.
The std-based weighting achieves this by discounting downstream tasks where all teams perform similarly, in analogy to:
A football match is a draw regardless if the end result is 1-1 or 8-8 — although the number of goals can have a marginal effect in a tournament.
We assign more importance to downstream tasks where participants score high AND when they distinguish themselves from the rest. More formally speaking: For a weight $w_{t}=\delta_{t}/\sum_{\tau}\delta_{\tau}$ with $\delta_{t}=\text{std}_{p}Q^{(p)}_{t}$ and the commonly accepted definition of variance
$(\text{std}_{p}A_{p})^{2}=\langle A_{p}^{2}\rangle_{p}-\bar{A}^{2}=\langle(A_{p}-\bar{A})^{2}\rangle_{p}$ (5)
where
$\langle f(X_{p})\rangle_{p}=\frac{1}{P+1}\sum_{p=0}^{P}f(X_{p})\quad\text{and}\quad\bar{A}=\langle A_{p}\rangle_{p}$ (6)
such that
$A_{p}=\bar{A}\quad\Leftrightarrow\quad\text{std}_{p}A_{p}=0~,$ (7)
the case $\delta_{t}\to 0$ for all $t$ may generate a numerical instability. However, our two distinct competition baselines
  • $p=1$: simple data aggregation of SSL4EO-S12 data cubes termed Baseline mean embeddings in the 2025 CVPR EARTHVISION data challenge with leaderboard mean Q-score $\langle Q_{t}^{(1)}\rangle_{t}=$-0.786
  • $p=0$: random embeddings termed Baseline random embeddings in the 2025 CVPR EARTHVISION data challenge with leaderboard mean Q-score $\langle Q_{t}^{(0)}\rangle_{t}=$-7.092
prevent $\delta_{t}=0$ in practice as verified by running the CVPR EARHTVISION data challenge over a month with more than 400 submissions from over 20 teams.
From a theoretical perspective, one may want to stabilize $w_{t}$ by adding a ghost task $t=0$ with variance $0<\delta_{0}=\epsilon\ll 1$ such that
$\delta_{0}=\sqrt{\left\langle Q_{0}^{(p)2}\right\rangle_{p}}>0\quad\text{setting}\quad\bar{Q}_{0}=0\quad\text{and defining}\quad R_{0}^{(p)}=0~.$ (8)
Abbreviating $\sum=\sum_{t}\delta_{t}$ we distinguish the cases
  • $\sum\gg\epsilon$: where $w_{0}=\epsilon/\sum\ll 1$ and $w_{t}=\delta_{t}/\sum$ leaving $s^{(p)}$ of Eq.˜4 intact to 0th order in $\epsilon$
  • $\sum\approx\epsilon$: where $w_{0}\approx\frac{1}{2}$ and $w_{t}\approx\frac{1}{2}\delta_{t}/\sum$ such that with $R_{0}^{(p)}=0$ the score $s^{(p)}$ in Eq.˜4 receives a discount factor $\tfrac{1}{2}$ which further increases for $\sum\to 0$ where $w_{0}\to 1$
For the CVPR EARTHVISION data challenge we ran NeuCo-Bench with task weighting with $1=\sum_{t}w_{t}$. When users simply want to benchmark their neural compression methodologies on a (sub)set of downstream tasks with known complexity without competing against other teams, the unweighted averaging is the preferred mode of operation for NeuCo-Bench. In Section˜A.3 we report on operational insights related to task weighting as observed in the context of the CVPR EARTHVISION data challenge. The results underline that the concept of task competitiveness bears further opportunities for continued research.
A.2 Platform and Infrastructure
Figure 6: Components and interaction of thedata challenge. The community platform Eval.AI (a) interacts with a virtual machine hosted at a cloud service (b). The virtual machine returns the quality score of the submission to the Eval.AI leaderboard and pushes updates to the custom leaderboard (c).
The core evaluation pipeline was implemented on a virtual machine (VM) with specifications:
  • Operating System/OpenStack Image: Ubuntu Jammy 22.04 LTS
  • CPU: 8 vCPUs, no GPU
  • RAM: 16GB
  • Disk: 20GB (OS) + 200GB (data storage)
running on top of the Juelich Supercomputing Center’s (JSC) OpenStack [^5] https://www.openstack.org cloud environment. The communication with the Eval.AI API for fetching submission data and writing results back to the Eval.AI leaderboard was based on the Eval.AI GitHub remote challenge evaluation template [^6] https://github.com/Cloud-CV/EvalAI-Starters/commit/8338085c6335487332f5b57cf7182201b8499aad utilizing the requests Python library. [^7] https://docs.python-requests.org Figure˜6 illustrates the entire setup: (a) the Eval.AI web interface [^8] https://eval.ai/web/challenges/challenge-page/2465 and a supplementary repository [^9] https://github.com/DLR-MF-DAS/embed2scale-challenge-supplement on one end, and (b) the evaluation procedure which runs on the VM at JSC, on the other end.
As evaluation method (Evaluate.py), NeuCo-Bench [^10] https://github.com/embed2scale/NeuCo-Bench/tree/040ca567da4d231ced78a16448d2039a0e871276 was incorporated into the Eval.AI remote challenge evaluation template running on the VM, cf. Local Repository – Evaluation Codebase in Fig.˜6. Updates to the Eval.AI challenge web interface got triggered by GitHub Actions. [^11] https://docs.github.com/en/actions In addition, the Supplementary Repository [^12] https://github.com/DLR-MF-DAS/embed2scale-challenge-supplement serves two purposes:
  • for the challenge participants to provide instructions and code examples with options to raise issues, and
  • to host a Custom Leaderboard implementing the global ranking introduced in Section˜A.1, not natively supported by Eval.AI
The VM runs a cronjob to restart Evaluate.py in case the application terminated. In fact, every minute NeuCo-Bench polls Eval.AI for new Submissions to score. Thereafter, the VM reports $Q$ of the evaluated submission to the Eval.AI Leaderboard. It also updates the global Custom Leaderboard in the GitHub Supplementary Repository.
A.3 Competition Analysis
Figure 7: The evolution of participant rankings in the challenge test phase. Lines correspond to participating teams and dots to submissions by the corresponding team. Team 10 is the simple mean baseline described in Section˜B.3.1, and Team 11 is a randomized baseline with randomly sampled, normally distributed embeddings.
The interaction between participants and organizers through GitHub issues allowed for transparent and traceable communication. In particular, we highlight an update to the challenge that improved comparability between participants by reducing variability in case the same submission is submitted multiple times. [^13] https://github.com/DLR-MF-DAS/embed2scale-challenge-supplement/issues/8
Other learnings from the develploment phase are:
  • Normalization of the target labels across all downstream tasks may be necessary to avoid hyperparameter tuning of the linear probe.
  • Before normalizing the target labels, the range of the target labels heavily affected the ability of the linear probe to learn a specific task within the given network initialization, learning rate and number of epochs.
In total, nine teams participated publicly in the final phase of the 2025 CVPR EARTHVISION data challenge, competing over scoring top rank and highest mean $Q$ value across all tasks.
In general, the ranking and the mean $Q$ value are close to identical. However, the team achieving third place upended the order of first and second place, with the effect that the runner-up team achieved a slightly higher mean $Q$-score than the winners. This effect is driven by a change in task weights caused by the third-place team’s performance. We note the dynamics around Submissions 14 and 15 as illustrated in Fig.˜7 where the ranking dynamics given a sequence of Submissions (dots) is documented: Team 1 through 9 are competing numerically indexed by their final position in the challenge ranking. Team 10 represents the simple mean baseline case described in Section˜5.3 with additional details in Section˜B.3.1, and Team 11 is a randomized baseline submitting randomly sampled, normally distributed embeddings.
A hallmark of our dynamic (global) ranking $R^{(p)}$ can be observed as follows: At submission 14, the submission of Team 3 modified the task weights such that the position of Team 4 and 5 were swapped, even though Team 3 ranked below the two other teams. The same occurred at submission 15, where the submission of team 3 changed the task weights to the benefit of Team 1. The adaptations of the task weights were small compared to the weights of the four tasks with highest weights—on the order of a few percent of the task weights. This drastic effect on the rank of the first two positions is partly due to Team 1 and 2 being neck and neck, Team 1 winning with a weighted average rank 2.31 and Team 2 coming second with 2.44, even though Team 2 scored 15.2 mean $Q$, ahead of Team 1 on 14.9.
The proposed task weighting method as defined by Eq.˜4 achieved to balance the importance of tasks. We noted that the agriculture and forest related tasks were well solved by several teams. The other tasks turned out more challenging. As expected, the (random) baselines were indicated low performers according to $R^{(p)}$. A limitation of the weighting we observed: Since the weighting of the tasks $t$ in Eq.˜4 is based on the variations of the participants for that given task $t$, a participant $p$ very poorly performing by design—such as the random baseline (Team 11)—artificially inflates the task weight when all the other participants perform well. A sensible extension of the NeuCo-Bench framework as discussed in Appendix˜B will depend on a careful design of downstream tasks and corresponding baselines.
Figure 8: Visualisation of downstream-task labels (top row) and corresponding Sentinel-2 images (bottom row). The biomass labels have been sourced from Karaman et al. (2025). Although Sentinel-1 is included in every data cube, it was excluded from this visualisation. The cloud and heat island labels are based on aggregated images.
Since the competition concluded in early April 2025, the one winner and the fourth-place team have open-sourced their solutions: the team achieving highest mean $Q$ is available, [^14] https://github.com/KerekesDavid/embed2scale-solution and the fourth-place team, who avoided (pre-)training completely by utilizing MOSAIKS (Rolf et al., 2021) is available, [^15] https://github.com/isaaccorley/temporal-mosaiks too.
Appendix B An Extendable Framework
Table 2: Qualitative comparison of our benchmark, PANGAEA, and GEO-Bench
ours PANGAEA GEO-Bench
Domain General purpose compression Geo. foundation models Geo. foundation models
Compute commodity hardware AI accelerator AI accelerator
Model access Not required Intermediate features Backbone finetuning
Tasks Classification Classification Classification
Regression Regression -
- Segmentation (focus) Segmentation
Leaderboard API JSON JSON -
Based on our insights from the CVPR EARTHVISION 2025 data challenge, we took our approach to the next level with the intention to build a community around benchmarking neural compression. Table˜2 provides a high-level comparison on how NeuCo-Bench fits into existing, popular geospatial benchmarking frameworks. In summary, NeuCo-Bench fills the following gaps:
  • Quantifies the quality of small embeddings based on a variety of downstream tasks without fine-tuning of any neural network backbone.
  • Provides a standalone toolkit for rapidly benchmarking any compressed embeddings beyond foundation models. In contrast to GEO-Bench and PANGAEA, the NeuCo-Bench framework is readily adapted to any compression scenario given:
    • Users provide embeddings $z$ where their encoder $E$ takes care of data formats.
    • Downstream tasks are shared with NeuCo-Bench as simple CSV files.
  • Supplies a multi-task performance metric that quantifies embedding size ($N$) vs. downstream accuracy ($Q$).
It is worth iterating that our dataset SSL4EO-S12-downstream dataset builds on the SSL4EO-S12 data structures part of GEO-Bench—thus, serving as a potential interface regarding synergies. We are currently forming the Earth2Vec community around NeuCo-Bench where more than 30 organizations from academia, the corporate world, and governments have joined.
B.1 Benchmark Tasks
Figure˜8 illustrates examples of Sentinel-2 inputs alongside the corresponding labels. Table˜3 lists a detailed overview of the derived 11 tasks, whereby 9 of these were used in the CVPR EARTHVISION 2025 data challenge and are highlighted with green check marks. Each task is linked to a HuggingFace identifier, following the definition in SSL (2025) and listed in Table˜3. Clouds and Nodata were not included in the CVPR competition but are provided with the release of NeuCo-Bench. The task Random provides randomly generated labels and associated data cubes from the Cloud task, which introduced an additional quality assessment.
The processing pipeline for all presented datasets utilized GEE GEE to download data cubes, applying a maximum cloud coverage filter of 10%, as provided by the GEE property CLOUD_COVER, except for the clouds task where no restrictions on cloud cover were enforced. Each data cube was aligned to the center of the corresponding label and processed to a size of 264 x 264 pixels. All sample locations with less than 4 images were discarded, following the requirements in Blumenstiel et al. (2025). Whenever possible, only locations that cover all four seasons for Sentinel-1 and Sentinel-2 were chosen. In case of missing latitude and longitude, locations were randomly selected from shapefiles representing regions such as mainland Europe or areas within the US Corn Belt. Comparison between the presented dataset and the data provided in Blumenstiel et al. (2025)show one major difference. In January 2022, ESA introduced a new baseline for Sentinel-2 data, effectively shifting all pixel intensities by 1000 units upward. The dataset presented in this work follows the format of GEE, i.e. removing this translation such that the minimum value for Sentinel-2 pixels are 0 both before and after the change by ESA. Blumenstiel et al. (2025) adheres to the ESA standard and enforces a lower bound for Sentinel-2 pixels of 1000, before and after the change. To allow for seamless integration between the two datasets, the dataloaders provided in NeuCo-Bench includes a setting that toggles a shift by 1000, aligning the distributions of the two datasets.
The Heat-island task required additional pre-processing, as Landsat-8 band 10 (B10) was utilized for label generation. This dataset considers only cities with populations exceeding 20,000 and a latitude between 8 and 70 north. The labels are based on all available Landsat-8 observations from June to the end of August for the years 2021 to 2024 inclusive. In addition, to reduce the impact of remaining clouds, any pixel with a combined brightness (red channel + blue channel + green channel) exceeding 30% of the maximum possible value or with a B10 temperature lower than 273 K are removed. Images with more than 10% removed pixels were dropped. The northernmost locations were verified to have average summer temperatures above freezing. For each location, the remaining images are flattened and concatenated over time, and then the mean and standard deviation are calculated from all pixels. The task is to estimate these spatio-temporal statistics.
Table 3: Descriptions of the downstream tasks provided by the initial release of our benchmark. The tasks used in thedata challenge are indicated with green check marks in the “Challenge” column. The task names correspond to the identifiers as used in the corresponding dataset released.
Task HuggingFace file Challenge Description
Biomass biomass_mean__regr Regression tasks: Biomass density (Mg/ha) mean and standard deviation estimated for pixel-level labels derived from GEDI.
biomass_std__regr
Crops crops__regr Regression task: Combined fraction of Soybean and Corn in the label image.
Landcover landcover_agriculture__regr Regression task: Percentage of agriculture pixels in the Corine Land cover image.
landcover_forest__regr Regression task: Percentage of forest pixels in the Corine Land cover image.
Clouds clouds_reg__regr Regression task: Average cloud cover fraction across four seasons in one year.
Heatisland heatisland_mean__regr Regression tasks: Summer surface temperature mean and standard deviation in Kelvin.
heatisland_std__regr
Nodata nodata__regr Regression task: Fraction of pixels with value zero in a Sentinel-2 image (based on all 13,260 available samples).
Random random_reg__regr Regression task: Random task with a majority of zero labels. The data cubes are the same as for Clouds.
random_cls__cls Classification task: Random binary classification for a majority of zero labels. The data cubes are the same as for Clouds.
Public vs. Secret Downstream Tasks. We released the hidden NeuCo-Bench downstream tasks (cf. Table˜3 with green check mark) after the conclusion of the 2025 CVPR EARTHVISION workshop to make publicly accessible the standalone SSL4EO-S12-downstream dataset for reasons of transparency, and to be used and contributed to by the neural compression community. As common with public benchmarks designed for standalone usage, we assume that NeuCo-Bench users would not jeopardize the developments of their own compressor $E$ by willingfully exploiting knowledge of the downstream tasks they test on. Removing and adding (mix-and-match) downstream tasks in NeuCo-Bench for a new competition avoids overfitting of a state-of-the-art compressor $E$. The process is as straightforward as uploading such data $x$ to HuggingFace, i.e.,
  1. 1.
    public: Each data point $x_{i}$ just needs a unique (identified by hash $i$) name (cf. e.g., directory data/ of SSL4EO-S12-downstream dataset) to upload corresponding …
  2. 2.
    secret till conclusion of competition: …label CSV files (cf. e.g., directory labels/ of SSL4EO-S12-downstream dataset, id column of CSV file) …
…for the NeuCo-Bench engine to perform its downstream task evaluations.
B.2 Standalone Python Implementation
In order for a clean separation of code from the open-source platform Eval.AI, we developed a minimal viable standalone Python code base [^16] https://github.com/embed2scale/NeuCo-Bench/tree/040ca567da4d231ced78a16448d2039a0e871276 to serve as plug-and-play for any larger ecosystem integrating of the core NeuCo-Bench framework. In fact, as Fig.˜6 demonstrates, the scoring for the Eval.AI leaderboard is entirely taken care of by NeuCo-Bench. Correspondingly, our framework commits an additional, customized leaderboard that globally depends on all submissions to a dedicated GitHub repository. [^17] https://github.com/DLR-MF-DAS/embed2scale-challenge-supplement/tree/272c911e8fb527d265de9be18fc81e16208ca0b6?tab=readme-ov-file#leaderboard
The NeuCo-Bench core evaluation.py functionality [^18] https://github.com/embed2scale/NeuCo-Bench/blob/040ca567da4d231ced78a16448d2039a0e871276/benchmark/evaluation/evaluation.py separately fetches
  • the user’s embeddings (submission), /path/to/submission_file.csv, and
  • the downstream task annotation data (labels), /path/to/annotation_directory/
as ASCII-formatted CSV files given predefined local paths and directories [^19] https://github.com/embed2scale/NeuCo-Bench/blob/040ca567da4d231ced78a16448d2039a0e871276/benchmark/main.py#L14-L19 as simple interface entirely independent of Eval.AI. Given any ranking procedure implemented, [^20] https://github.com/embed2scale/NeuCo-Bench/blob/040ca567da4d231ced78a16448d2039a0e871276/benchmark/evaluation/results.py#L43-L80 the resulting leaderboard is saved as human-readable JSON file [^21] https://github.com/embed2scale/NeuCo-Bench/blob/040ca567da4d231ced78a16448d2039a0e871276/benchmark/evaluation/results.py#L8-L22 in a corresponding /path/to/results_directory/. For downstream (binary) classification tasks, the confusion matrix and related scores such as precision, recall, F1, and overall accuracy are calculated along with the ROC-AUC-score (area under Receiver-Operator-Characteristic graph). For regression, the R-squared, mean squared, and mean absolute errors are computed. [^22] https://github.com/embed2scale/NeuCo-Bench/blob/040ca567da4d231ced78a16448d2039a0e871276/benchmark/evaluation/metrics.py
To serve as seed towards an open-source and open science community, we designed the standalone Python implementation of NeuCo-Bench modular for easy extension. Depending on compute resources, we encourage future contributions to add novel probing models, cross validation schemes, and performance scores (cf. Eq.˜3) beyond the current. [^23] https://github.com/embed2scale/NeuCo-Bench/blob/040ca567da4d231ced78a16448d2039a0e871276/benchmark/evaluation/linear_probing.py As a bonus, our standalone implementation allows to store plots of loss curves, linear correlation of regression tasks, and a confusion matrix for classification on disk. [^24] https://github.com/embed2scale/NeuCo-Bench/blob/040ca567da4d231ced78a16448d2039a0e871276/benchmark/evaluation/visualisations.py
Running NeuCo-Bench standalone on the command line reduces to something as simple as:
1python main.py \
2 --annotation_path /path/to/annotation_directory/ \
3 --submission_file /path/to/submission_file.csv \
4 --output_dir /path/to/results_directory/ \
5 --config /path/to/config.yaml \
6 --method_name ’your-method-name’ \
7 --phase ’phase-name’
where your-method-name and phase-name are free strings to define an output (sub-)directory
/path/to/results_directory/phase-name/your-method-name_YYYYMMDD_HHmmSS
with YYYYMMDD a date of year YYYY and zero-padded numerical month MM and day DD. HHmmSS indicates a time of the day in hours HH, minutes mm, and seconds SS, accordingly. A YAML file /path/to/config.yaml specifies details of the evaluation such as:
1embedding_dim: 1024 # number of embedding dimensions
2batch_size: 64 # batch size for (linear) probing
3epochs: 20 # number of epochs to optimize the (linear) probe for
4learning_rate: 0.001 # learning rate to optimize with
5k_folds: 40 # number of cross-validations to generate statistics over
6standardize_embeddings: true # standardize embeddings by their global mean and std
7normalize_labels: true # normalize in range [0,1]
8task_filter: false # all in /path/to/annotation_directory/ per default
9 # example: ["biomass_mean", "biomass_std"]
10# etc.
B.3 General Evaluation Setup
We expand on the methods introduced in Section˜5.3, applying off-the-shelf embedding methods to the downstream tasks defined in Section˜B.1. Unlike the CVPR challenge, where data were partitioned into development and evaluation phases, results here are computed on the entire downstream dataset. Unless otherwise stated, evaluation follows the challenge protocol and used hyperparameters: $E=20$ epochs, $k=50$ training and test set splits, and a learning rate of $10^{-4}$. We report raw $R^{2}$ scores, clipping negatives to $[0,1]$ for figures.
For the results in Section B.3.1, we use embeddings of size $N=1024$, consistent with the challenge setup. In Section B.3.3, we relax this requirement to investigate the effect of varying embedding dimensionality and explore decoder choices beyond the linear probes proposed as default.
Figure 9: Illustration of pre-encoding vs. post-encoding aggregation. In post-encoding, each seasonal image is encoded separately before combining embeddings, which mitigates outlier effects (e.g., snow) but increases runtime fourfold.
B.3.1 Method Details
Averaging Baseline. As a simple informative reference, we construct a Mean baseline by strongly downsampling and averaging the SSL4EO-S12 data cubes. First, we reduce the spatial resolution of each of the 27 channels from 264x264 pixels to 8x8 by bi-linear interpolation. Next, we exploit correlation as visualized by Fig.˜10 reducing the number of channels from 27 down to four. We average the channels B1 through B9 of both S2L1C and S2L2A and we do similar for channels B11 and B12, respectively. Channel B10 of S2L1C is kept separate since no corresponding band exists in S2L2A. Seasonal snapshots are kept separate, yielding $\times8$$\times$$\times4$ values, flattened to $N=1024$. This baseline indicates how much task-relevant information survives coarse spatial and spectral aggregation.
Figure 10: Pearson correlation coefficient matrix of the 27 data cube channels. In here, we abbreviate the channels of the Sentinel-2 L1C and L2A products as L1C and L2A, respectively appending the channel name (B1, B2, …).
Neural rate–distortion compressors. We adopt the Factorized Prior autoencoder of Ballé et al. (2016) for Sentinel-2 L1C imagery. Models use 256 intermediate channels and 128 latent channels, trained with loss $\mathcal{L}=R+\lambda D$, where $D$ is MSE distortion and $\lambda\in\{0.025,0.1,0.5\}$. These compressors match or exceed JPEG2000 PSNR at roughly half the bitrate (Fig. 11). At inference, we pool latents to $4\times 4$, flatten to 2048, and average adjacent channels to yield 1024-dim embeddings.
Self-supervised foundation models (FMs). We evaluate unimodal FMs pretrained on Sentinel-2 L1C (Wang et al., 2023b), as well as multimodal FMs (DOFA (Xiong et al., 2024), TerraMind (Jakubik et al., 2025)). CNN backbones are downsampled by spatial averaging to 2048 dims, then halved to 1024. ViT-B/16 backbones average patch embeddings, padding from 768 to 1024 dims.
Temporal aggregation. Each sample contains four seasonal snapshots. In pre-encoding, channels are averaged across seasons before encoding. In post-encoding, each snapshot is encoded separately and averaged. As shown in Fig. 9, post-encoding better handles seasonal outliers (e.g., snow), at the cost of increased compute.
B.3.2 Key Results Across Embedding Methods
Figure 11: Rate–distortion performance of the Factorized Prior neural compressor, demonstrating superior compression quality over the JPEG 2000 baseline.
Table˜4 summarizes linear-probe $R^{2}$ scores with pre- and post-encoding aggregation.
Temporal aggregation. Post‐encoding aggregation yields absolute $R^{2}$ gains for almost every method and task. For instance, the TerraMind cloud performance increases from 0.434 to 0.731 (+0.297), and landcover tasks by +0.011–0.016.
Neural compressors. Reducing $\lambda$ from 0.1 to 0.025 yields small average improvements, whereas $\lambda=0.5$ degrades performance and produces negative $R^{2}$ on several tasks.
ViT backbones. MAE achieves balanced performance across tasks, while contrastive DINO and MoCo excel on semantic landcover tasks (e.g., DINO post‐encoding achieves $R^{2}=0.863$ on crop fraction) but are less competitive on geophysical regression.
CNN backbones. Although CNN variants perform strong on landcover tasks, they fall behind ViTs on geophysical regressions. They benefit from post‐encoding aggregation—for instance, DINO CNN’s biomass mean improves from negative scores to 0.117.
Multimodal models. TerraMind consistently outperforms all other embeddings, achieving the highest $R^{2}$ across tasks. DOFA, while generally scoring below TerraMind, still achieves consistent performance across all task.
Table 4: Full per‐task $R^{2}$ scores for all tested embedding methods. For all methods, we report scores for Pre-Encoding Temporal Aggregation (Pre-Enc) and Post-Encoding Temporal Aggregation (Post-Enc). Methods are grouped into the main categories: Multimodal (Terramind, DOFA), ViT FMs, CNN FMs, Neural Compressors, and Averaging Baseline.
Method Biomass Mean Biomass Std Crops Clouds Landcover Agriculture Landcover Forest Heatisland Mean Heatisland Std
Averaging Baseline -0.552 -0.426 -0.061 -2.293 0.126 0.216 -0.962 -1.157
Neural Compressors (Factorized Prior)
$\lambda$ = 0.025 (Pre-Enc) 0.186 0.132 0.423 0.244 0.463 0.246 0.321 -0.114
$\lambda$ = 0.025 (Post-Enc) 0.195 0.140 0.338 0.288 0.449 0.256 0.315 -0.113
$\lambda$ = 0.1 (Pre-Enc) 0.127 0.078 0.388 0.222 0.447 0.278 0.238 -0.246
$\lambda$ = 0.1 (Post-Enc) 0.129 0.078 0.357 0.266 0.464 0.333 0.219 -0.270
$\lambda$ = 0.5 (Pre-Enc) -0.037 -0.062 0.364 0.129 0.464 0.319 0.067 -0.563
$\lambda$ = 0.5 (Post-Enc) -0.036 -0.071 0.325 0.200 0.478 0.343 0.033 -0.628
CNN FMs (Unimodal)
DINO (Pre-Enc) -0.031 0.015 0.749 -0.171 0.855 0.838 0.396 -0.334
DINO (Post-Enc) 0.117 0.088 0.826 0.147 0.879 0.865 0.483 -0.221
MoCo (Pre-Enc) -0.164 -0.111 0.688 -0.228 0.835 0.808 0.236 -0.502
MoCo (Post-Enc) -0.048 -0.030 0.780 0.216 0.864 0.844 0.345 -0.334
ViT FMs (Unimodal)
DINO (Pre-Enc) 0.159 0.157 0.755 0.061 0.822 0.812 0.446 -0.169
MoCo (Pre-Enc) 0.294 0.229 0.706 0.242 0.790 0.778 0.477 0.113
MoCo (Post-Enc) 0.338 0.259 0.751 0.409 0.825 0.814 0.506 0.133
Multimodal FMs
TerraMind (Pre-Enc) 0.492 0.366 0.836 0.434 0.907 0.892 0.676 0.193
TerraMind (Post-Enc) 0.528 0.390 0.879 0.731 0.918 0.908 0.691 0.226
DOFA (Pre-Enc) 0.311 0.234 0.622 0.409 0.762 0.747 0.521 0.156
DOFA (Post-Enc) 0.373 0.269 0.587 0.518 0.777 0.773 0.560 0.207
Intermediate Layers. To probe how different layers capture task‐relevant features, we extracted intermediate embeddings from the contrastive ViTs. As shown in Fig.˜12, low‐level regression tasks (biomass mean/std, heat‐island mean/std, cloud fraction) peak at shallow layers (layer 1), whereas higher‐order semantic tasks (crop fraction, landcover proportion) reach their best performance at deeper layers.
Figure 12: Layer-wise $R^{2}$ performance of ViT-based foundation models across downstream tasks. Shallow layers capture low-level signals useful for regression tasks (e.g., biomass, heat-islands, clouds), while deeper layers improve performance on semantic classification tasks (e.g., crops, landcover).
B.3.3 Ablations and Extended Analyses
Decoder Analysis (Linear vs. Non-Linear Probes).
As part of our evaluation design, we explored the impact of decoder complexity on downstream task performance. While linear probing is the default protocol, we deliberately investigated non-linear alternatives to assess whether additional decoder capacity meaningfully improves results.
Figures 5(b) and 13 compare linear probes with one- and two-hidden-layer MLP decoders. We observe three consistent findings:
  • Stable rankings. The relative ranking of embedding methods remains nearly the same across probe types, indicating that differences between methods are not an artifact of probe capacity.
  • Marginal gains for strong embeddings. Top-performing embeddings (e.g., TerraMind, MAE) improve by less than 0.06 $R^{2}$ on average when switching to non-linear probes, demonstrating that these embeddings are already highly linearly expressive.
[…]
Interestingly, weaker embeddings benefit disproportionately from non-linear probes, suggesting that added decoder complexity can compensate for lower-quality representations. However, this comes at substantial computational cost.
Taken together, these results highlight that linear probing is not only efficient but also a discriminative evaluation strategy: it faithfully reflects the intrinsic quality of embeddings while enabling scalable benchmarking. Non-linear decoders may be useful for future extensions to more complex tasks (e.g., pixel-wise segmentation), but for the image-level tasks studied here, linear probing provides a robust and interpretable measure of embedding quality.
Figure 13: Per-task results for linear vs. non-linear probes. Non-linear decoders benefit weaker embeddings but have little effect on top-performing methods.
Figure 14: Embedding size ablation for ViT-based models. Performance peaks at the native patch embedding size and drops with reduced dimensions.
Figure 15: Embedding size ablation for CNN-based models. Optimal performance occurs between 128–1024 dimensions, with degradation outside this range.
Embedding Size Ablations.
Figures 14 and 15 show ablation results on embedding dimensionality for ViT-based and CNN-based models, respectively. For CNN backbones, performance generally peaks in the range of 128–1024 dimensions, with larger or smaller embeddings leading to consistent performance drops. ViT-based embeddings, by contrast, are most effective at their natural patch-token dimension, and reductions tend to degrade task performance. Notably, the benefit of larger embeddings is limited: increases beyond 1024 dimensions yield negligible accuracy improvements while substantially raising computational demands and probe parameter counts. These results justify the use of 1024-dimensional embeddings as a balanced default in the run data challenge, while also illustrating NeuCo-Bench’s flexibility for exploring embedding-size vs. utility trade-offs in future studies.
B.4 Compute Resources & Tuning Parameters
Table 5: Empirical runtime (in seconds) for different tasks under varying embedding size (NN), number of epochs ($E$), and number of CV folds ($k$) on single-CPU commodity hardware. Vertical lines separate configurations with different embedding sizes.
Task (# samples) $N=1024$ $N=512$ $N=2048$
$E=10$ $k=40$ $E=20$ $k=20$ $E=20$ $k=40$ $E=40$ $k=40$ $E=20$ $k=80$ $E=20$ $k=20$ $E=20$ $k=40$ $E=20$ $k=20$ $E=20$ $k=40$
Biomass (2415) 4.23 4.24 7.98 16.01 16.54 3.87 7.54 4.25 8.32
Crops (3355) 5.60 5.56 11.26 22.15 22.00 5.09 10.35 5.94 12.34
Clouds (1140) 2.04 2.03 3.94 7.76 7.59 1.91 3.69 2.13 4.12
Landcover Agriculture (4691) 7.70 8.06 15.75 31.03 30.75 7.08 14.26 8.07 16.40
Landcover Forest (4691) 7.86 7.81 15.48 31.06 30.47 7.08 14.37 8.06 16.64
Heatisland (1659) 2.79 2.80 5.53 10.84 10.83 2.66 5.25 2.89 5.85
No-data (13260) 22.45 22.45 44.30 88.68 88.18 20.46 42.03 23.18 46.60
At the 2025 CVPR EARTHVISION data challenge, NeuCo-Bench completed evaluations for a single submission within about 10 minutes for the embedding dimension of $N=1024$, the number of epochs per fold are $E=20$, the number of training- and test set splits had been set to $k=40$ in the development phase and $k=200$ in the evaluation phase. The evaluation script ran across a diverse set of eight downstream tasks for real-world geospatial applications. As alluded in Section˜B.2, users can flexibly adjust evaluation parameters in order to tune the runtime of the standalone implementation of NeuCo-Bench:
  • Embedding dimension ($N$, embedding_dim)
  • Number of epochs per CV fold ($E$, epochs)
  • Number of CV folds ($k$, k_folds)
  • Choice of tasks included (task_filter)
Empirical runtime measurements confirm an approximately linear scaling w.r.t. both, the number of epochs $E$ and the number of cross-validation folds $k$. A similar scaling behavior was numerically verified for the dataset size (# samples) for fixed downstream task. Runtimes are further influenced–—though to lesser extent—–by the embedding dimensionality $N$. For example, increasing the embedding size from 512 to 1024 dimensions results in a runtime increase of approximately 5% to 10% across tasks. For $N=1024$ to $N=2048$ dimensions implies an additional increase of about 5% to 15%—depending on the task dataset size. Such a sub-linear scaling may be attributed to computation overheads and system-level inefficiencies. Those may reduce the relative computational costs when increasing the embedding dimensionality $N$. Table˜5 lists a collection of recorded execution times (in seconds) for various parameter configurations per downstream task. All runtimes were measured on a single commodity ARM64 CPU with 16 cores (4.06 GHz) and 64 GB of RAM.
B.5 Licenses for Data & Software
NeuCo-Bench builds on open-source software and is released under the Apache 2.0 license publicly available at https://github.com/embed2scale/NeuCo-Bench. All package dependencies are listed in the requirements.txt [^25] https://github.com/embed2scale/NeuCo-Bench/blob/main/requirements.txt file, and those are licensed under widely-accepted open-source terms [^26] https://opensource.org/licenses, including BSD, MIT, PSFL, The Unlicense272727code-equivalent to CC0 data licenses, MPL-2.0282828weak copy-left that allows for integration with non-copyleft licenses, and Apache. These permissive licenses allow academic research and commercial use, making them fully compatible with the chosen Apache 2.0 license. Table˜6 lists all data currently included in NeuCo-Bench along with their origin. Google Earth Engine (GEE) (GEE, ) was utilized as the primary platform for downloading downstream task data as introduced in Section˜B.1. Future data and code contributions to NeuCo-Bench are required to be licensed under CC-BY 4.0 and Apache 2.0, respectively.
Table 6: List of licenses related to datasets currently included in our benchmark. All of these, except Clouds, are available in GEE. Table˜1 lists years of target labels, ranging from 2018 through 2024.
Dataset Origin of Data License
Sentinel-1 & -2 ESA / Copernicus CC BY-SA 3.0 IGO
Landsat-8 USGS (Observation and Center, 2020) Public Domain
CDL USDA NASS Cropland Data Layers (Boryan et al., 2011) Public Domain
CORINE European Environment Agency (EEA), European Union Full, Open,
Copernicus Land Monitoring Service (European Environment Agency (2018), EEA) and Free Access
CloudSen12+ CloudSEN12 project (Aybar et al., 2024) CC0 1.0
GEDI NASA (Dubayah et al., 2022) Public Domain
We note that the current NeuCo-Bench implementation lists CUDA packages covered by a proprietary NVIDIA license [^29] https://docs.nvidia.com/cuda/eula/index.html. However, we do neither bundle nor redistributes corresponding binaries. Users and contributors to NeuCo-Bench that share related docker containers need to explicitly attribute NVIDIA’s license. Fortunately, and as alluded in Table˜5 and Section˜A.2, NeuCo-Bench runs swiftly in a VM with commodity hardware specifications on CPU compute, only. Accordingly, the standalone NeuCo-Bench implementation introduced in Section˜B.2 can be started with (Bash) environment variable CUDA_VISIBLE_DEVICES=’’ to avoid usage of GPU resources.